OpenStudy (anonymous):

If tanhx = 12/13, find the values of the other hyperbolic functions at x.

OpenStudy (anonymous):

\[\tanh x = \frac{\sinh x}{\cosh x}\]

OpenStudy (anonymous):

\[\frac{ 1 }{ coshx } = \sqrt{1-\tanh^2x}\]

OpenStudy (anonymous):

\[sinhx = \sqrt{\cosh^2x-1}\]

OpenStudy (anonymous):

Did I misunderstand the question?

OpenStudy (anonymous):

Then you plug in the values and now you have to find the reciprocals.

OpenStudy (anonymous):

Wait.

OpenStudy (anonymous):

\[\tanh x = \frac{\sinh x}{\cosh x} = \frac{12}{13}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

But that doesn't necessarily mean that \(\sinh(x)=12\) That only give you the proportion.

OpenStudy (anonymous):

So use \[ (13c)^2-(12c)^2=1 \]and solve for \(c\) I guess.

OpenStudy (anonymous):

I knew it was too good to be true. Carry on. :D

OpenStudy (anonymous):

It's too tedious

OpenStudy (anonymous):

Well consider if \(\tan(x)=2/5\). Clearly \(\sin(x)\neq 2\) since it never gets that high.

OpenStudy (anonymous):

Yeah coshx = 13/5

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%2813c%29%5E2%E2%88%92%2812c%29%5E2%3D1 This tells us \(c=\pm 1/5\) So:\[ \sinh(x)=\pm \frac {12}{5}\quad \cosh(x)=\pm \frac {13}{5} \]Both are positive or both are negative.

OpenStudy (anonymous):

Would have to know the quadrant to tell.

OpenStudy (anonymous):

So I should add the + or - sign ?

OpenStudy (anonymous):

The reciprocals should be easy then, yeah?

OpenStudy (anonymous):

Make it only positive, then say another possibility is that both are negative.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

Or you could be really cute and say \[ |\cosh (x)|=\frac {13}{5} \]

OpenStudy (anonymous):

hahahaha!

OpenStudy (anonymous):

Technically you can't get any more accurate.

OpenStudy (anonymous):

Wait how did you figure out what coshx and sinhx were o_O

OpenStudy (anonymous):

Started with:\[\tanh x = \frac{\sinh x}{\cosh x} = \frac{12}{13}\] So I knew\[\sinh x=12c\quad \cosh x=13c\]

OpenStudy (anonymous):

I used the property: \[ \cosh^2x-\sinh^2 x=1 \]To get\[ (13c)^2−(12c)^2=1 \]

OpenStudy (anonymous):

Found out that \(|c|=1/5\)

OpenStudy (anonymous):

Ah, wow your way is much better than mine, and much simpler.